Four kids took a math test and gathered an average of a 75. Jenny had a score 10 more than Alex and 25 more than Nick. Lynn had a score of 86 and was the highest score out of the four. What is Jenny's score?

Found 2 solutions by Boreal, math_helper:
Answer by Boreal(15235)

The total score for all 4 is 300.
Lynn had 86
The other 3 had a combined score of 214
Nick is x
Alex is x+15
Jenny is x+25
Therefore, 3x+40=214
3x=174
x=58
Jenny got an 83 ANSWER
check--scores were 86,83,73,58

Answer by math_helper(2461)


Let J, A, N, L = scores for each person
(1) (J+A+N+L)/4 = 75 —> J+A+N+L=300
(2) J=A+10 —> A = J-10
(3) J=N+25 —> N = J-25
(4) L=86

Using rightmost equations, substitute into (1) for A, N, and L:
J+(J-10)+(J-25)+86 = 300

You can solve this for J and you are done.


I got J=83, A=73, N=58 (and with L=86, the average does check at 75)